Previous

Calendar

August

2008

Next
Sunday Monday Tuesday Wednesday Thursday Friday Saturday

27

JD209

28

JD210
Site Visit (SpB/JB, MP/SH/BH/NM, NT/SL): Details , Photos (11)

worked on PAR setup (worked on two trees in grove 2)

swapped 15 flasks at the west robot

Cool and overcast in morning--some rain and a bit of lightning/thunder in afternoon

29

JD211

30

JD212
High Ozone alerts issued for Denver-metro and Front Range region

31

JD213

1

JD214
Clear skies (and HOT) all day today

JB with Merced

2

JD215

3

JD216

4

JD217
Site Visit (SpB/JB, SH/SL): Details , Photos (17)

from 11:30-12:20, messed around with 16m and 25m prop-vanes (to determine the angle of vane relative to the boom).

at approx 12:30 MST, Changed profiler (li-6251) dessicant

from 12:44-12:50 MST, slight change to N2 pressure (from 20 to 18)

discovered slow ozone (TECO) pump was dead.

swapped 15 flasks at the west robot

Grey jeep had a flat tire (found this way in the parking lot); exchanged for teal jeep.

Big Cumulonimbus clouds up near the divide. Thought it was going to dump rain, but never really happened.

5

JD218
Site Visit (JB, MP/SH/JH, LL, NT/SL): Details

worked on PAR de-bugging

6

JD219

7

JD220

8

JD221
Site Visit (SpB/JB, SH/SL): Details , Photos (12)

at 10:00 MST, added 4 deg S-N tilt to KZ PAR sensor

from 10:00-11:30, messed around with 16m and 25m prop-vanes (using sextant to determine the angle of boom relative to N/E/S/W).

from 10:00-11:30, 15-m prop-vane held at constant angle

brought 1 new N2 to tower, brought 1 empty N2 back to Ramaley

span @ 275 psi, N2 @ 1000 psi

~10:00 AM MST, adjusted H20 zero on the LI-6262 from ~900 to ~ 840

JB with CU

9

JD222

10

JD223

11

JD224
Site Visit (SpB/JB, SL): Details , Photos (12)

at 11:32 MST, start running ops46, moved 2-d handar, from chan 200 (UC logger) to chan 202 (UCB logger)

from 11:07-11:30, tested out new "quacker" (Dell Inspiron 7500)

12

JD225

13

JD226

14

JD227

15

JD228
Wet and rainy (all day).

16

JD229
First snowfall of the season (accum on trees).

17

JD230

18

JD231 (mostly sunny, cooler)
SpB Off
Site Visit (JB, SL): Details , Photos (6)

brought two more N2 cylinders up to the site from Ramaley and stocked these at the tower

snapped new phenology pics

started a leak check in the trailer with the new co2 regulator (i.e. new stem and nipple)

swapped 15 flasks at the west robot

N2 @ 700 psi, span @ 150 psi

swapped span cylinders (retired cc199497, started cc282376 @ 1925 psi

swapped moisture trap and indicator at the LI-6251 (zero was at 0.100, so I re-zeroed it to 0.000)

19

JD232

20

JD233

21

JD234
Site Visit (JB--Merced work day)

checked the data system and discovered that LL had accidentally shut down russter2; successfully restarted russter2 and the data system

22

JD235

23

JD236

24

JD237

25

JD238
CU classes begin

Site Visit (JB, NT): Details

swapped 15 flasks

removed ATI 2-D sonic from the tower and brought to ATI in Longmont for test/repair

26

JD239 (windy)
Site Visit (JB, NT): Details

downloaded forestcam photos

downloaded from soil and Tcs

span @ 1850 psi, N2 @ 500 psi

removed KZ par sensor from the tower

cleaned krypton

changed out the magnesium perchlorate and indicator at the LI-6251

27

JD240

28

JD241

29

JD242

30

JD243

31

JD244

1

JD245
CU holiday (Labor Day)

2

JD246
Site Visit (JB, NT): Details

N2 @ 300 psi, span @ 1700 psi

swapped 15 flasks at the west robot

3

JD247

4

JD248
JB at site (Merced day)

span @ 1700 psi

swapped N2 tanks at the tower. retired N2 cc193097 and started cc1255575 @ 1925 psi

SpB Returns

5

JD249

6

JD250